40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 22 18:42:10 2011, kiwamu, Summary, Green Locking, Some issues on the Y end green PDH locking 
    Reply  Wed Nov 23 00:00:53 2011, Zach, Summary, Green Locking, Some issues on the Y end green PDH locking 
    Reply  Sat Nov 26 19:07:40 2011, kiwamu, Summary, Green Locking, AM trnasfer function of the Y end laser with doublin crystal AMTF_lightwave.pngAMTF_cailbrated.bod.zip
Message ID: 5980     Entry time: Tue Nov 22 18:42:10 2011     Reply to this: 5983   6018
Author: kiwamu 
Type: Summary 
Category: Green Locking 
Subject: Some issues on the Y end green PDH locking 

[Rana / Kiwamu]

 As a part of the ALS noise budgeting we took a look at the Y end PDH setup to see if we are limited by an effect from the Amplitude Modulation (AM).

Then we found two issues :
 (1) a big variation in AM transfer function from the laser PZT to the intensity of the frequency-doubled laser. We haven't figured out the reason yet,
 (2) some of the optics and their mounts need to be refined.

 


(AM transfer function)

 One of the suspicious noise source of the Y arm ALS was an AM effect in the Y end green PDH locking.
A possible scenario is that: there is some amount of the offset in the PDH signal due to the AM at the modulation frequency,
and it allows the intensity noise to couple to the laser frequency, which we want to suppress.
 So we wanted to check if the measured AM (#2799) at 1064 nm  is still true at 532 nm.
The problem right now is that : every time we measured the AM transfer function by exciting the laser PZT with swept sine,
the transfer function varied by 20 dB, with average response of 50 dB. And there was no repeatability.
We were using the PD which is for the green PDH signal and the single-bounced light from ETMY.
The measurement was done in a frequency band of 100 - 400 kHz where we expect a couple of sharp notches.
Perhaps we should try the same measurement with IR first to make sure we are doing a right thing, and then do it with the frequency-doubled laser.

 

(Y table setup needs more improvements)

  We found some optics and their mounts which need to be refined.
Here is a list which we briefly made at the time.
  • Use washers
  • Beam clipping in Green Faraday and the very last mirror
  • Use two screws and wide base plate
  • Tune PPKTP PID parameters
  • Remove flipper mirror
  • Move the mechanical shutter to where the beam size is smaller
  • Put a beam damp for the reflected light from the PD
  • Cable rack
  • Improve the incident angle on the last two launching mirrors
ELOG V3.1.3-